OpenStudy (anonymous):

Find the vertex to the quadratic equation: 3x^2-17x-6

OpenStudy (cwrw238):

= 3 (x^2 - (17/3) x ) - 6 = 3[ (x - (17/6)^2 - 289/36)] - 6 = 3[(x - 17/6)^2 ) - 289/12 - 6 = 3( x - 17/6)^2 - 24.08 - 6 = 3(x - 17/6)^2 - 30.08 vertex is at (17/6 , -30.08) = ( 2.83, -30.08)